# ALMA: a mathematics-driven approach for determining tuning parameters in generalized LASSO problems, with applications to MRI

### v1

@@ IV-E Comparison with the L-curve parameter
For a comparison, we report on the quantitative measurement regarding the reconstructions obtained utilizing the parameter $\lambda_{L}$, computed by $L-$curve. These raconstructions exhibit an average $mSSIM\geq 0.99$, across noise levels and undersampling rates (see fig. 11), they also exhibit an average pSNR of $42.6351$ and an average CJV of $0.035$. Therefore, they perform well regarding the metrics criterion. Moreover, the tuning parameter that minimizes the mSSIM ($\lambda_{mSSIM}$) is on average $0.6$ times the corresponding $\lambda_{L}$, i.e. $\lambda_{mSSIM}\approx 0.6\cdot\lambda_{L}$, while the tuning parameters optimizing the pSNR and the CJV ($\lambda_{pSNR}$ and $\lambda_{CJV}$, resp.) are on average $0.5$ times $\lambda_{L}$ (see fig. 11). All these data together show that the performance of the L-curve parameter is slightly better compared to the performance of the ALM. However, the difference is marginal and amounts to only a matter of decimals. Moreover, ALMA is an iterative procedure to compute tuning parameters and, consequently, solving TV-LASSO, whereas the L-curve method is heuristic and does not involve an iterative process. Despite yielding similar results, ALMA’s iterative nature ensures a more robust and accurate determination of tuning parameters for solving TV-LASSO, offering a superior alternative to the heuristic approach of the L-curve method. Therefore, both methods perform similarly overall, indicating that the ALM is a reliable parameter, performing on par with the well-established L-curve method.
@@
Figure 9: Means and standard deviations of the number of iterations, the three metrics, and the magnitude ratios across undersampling rate and noise level, for the reconstructions obtained using the ALM. Observe that the average number of iterations is 7, and the three metrics exhibit average values of approximately 0.9951, 42.2401 dB, and 0.0367, respectively, indicating optimal performance of ALMA.
V Discussion
@@
Figure 10: Example of reconstruction of simulated MRI data ($NL_{\%}=7/100$, $UR_{\%}=10/100$) with different tuning parameters: the ALM and the tuning parameter computed with the L-curve method, the tuning parameters that optimize each metric and tuning parameters that are powers of 10. The corresponding quantitative quality assessments are reported below each reconstruction. Observe that the quality of the reconstructions obtained using the ALM and the L-curve parameter are nearly indistinguishable, and close to the quality of the images reconstructed with the tuning parameters that optimize the metrics.
@@
We also compared the performance of ALMA against the well-established L-curve method. As presented in fig. 9 and 11, and exemplified in fig. 10, both methods show comparable performance across the metrics. However, ALMA stands out as the first iterative algorithm to compute the parameter dynamically during reconstruction. This active, iterative approach grants ALMA significant advantages in computational efficiency and ease of implementation, which are particularly beneficial in clinical settings where a rapid and optimized reconstruction is essential.
@@
A Refer to caption  B Refer to caption
Figure 12: Reconstructions of the Shepp-Logan brain phantom under radial sampling, with $NL_{\%}=15/100$ and $UR_{\%}=15\%$. The reconstruction obtained with the ALM is displayed on the left. The reconstruction obtained with the parameter that maximizes the mSSIM is displayed on the right. Observe that the two images are almost indistinguishable.
VI Conclusions and future directions

### v2

@@ 3.5 Comparison with the L-curve parameter
For a comparison, we report on the quantitative measurement regarding the reconstructions obtained utilizing the parameter $\lambda_{L}$, computed by $L-$curve. These raconstructions exhibit an average $mSSIM\geq 0.99$, across noise levels and undersampling rates (see figure 8), they also exhibit an average pSNR of $42.6351$ and an average CJV of $0.035$. Therefore, they perform well regarding the metrics criterion. Moreover, the tuning parameter that minimizes the mSSIM ($\lambda_{mSSIM}$) is on average $0.6$ times the corresponding $\lambda_{L}$, i.e. $\lambda_{mSSIM}\approx 0.6\cdot\lambda_{L}$, while the tuning parameters optimizing the pSNR and the CJV ($\lambda_{pSNR}$ and $\lambda_{CJV}$, resp.) are on average $0.5$ times $\lambda_{L}$ (see figure 8). All these data together show that the performance of the L-curve parameter is slightly better compared to the performance of the ALM. However, the difference is marginal and amounts to only a matter of decimals. Moreover, ALMA requires approximatively $7\pm 3$ reconstructions in average (since a reconstruction occurs at each iteration) and in any case less that 30 reconstructions in our experiment, while the L-curve method required 41 reconstruction to reach the result exhibited here for comparison. Finally, the L-curve method is equivalent to an exhaustive searche and fails in principle if the lagrange multiplier to approximate is out of the bound of the exaustive searche. In contrast, ALMA do not require any apriori knowledge about the lagrange multiplier to estimate in order to succeed. All these facts taken together can be interpreted as a superiority of ALMA.
@@
Figure 7: Means and standard deviations of the three metrics, and the magnitude ratios across undersampling rate and noise level, for the reconstructions obtained using the ALM. Observe that the average number of iterations is 7, and the three metrics exhibit average values of approximately 0.9951, 42.2401 dB, and 0.0367, respectively, indicating optimal performance of ALMA.
4 Discussion
@@
Figure 9: Example of reconstruction of simulated MRI data ($NL_{\%}=7/100$, $UR_{\%}=10/100$) with different tuning parameters: the ALM and the tuning parameter computed with the L-curve method, the tuning parameters that optimize each metric and tuning parameters that are powers of 10. The corresponding quantitative quality assessments are reported below each reconstruction. We observe that the quality of the reconstructions obtained using the ALM and the L-curve parameter are nearly indistinguishable, and close to the quality of the images reconstructed with the tuning parameters that optimize the metrics.
@@
We also compared the performance of ALMA against the well-established L-curve method. As presented in figures 7 and 8, and exemplified in fig. 9, both methods show comparable performance across the metrics. However, ALMA stands out as the first non-trained and self-calibrated algorithm that does not need any apriori knowledge on the lagrange multiplier to estimate and is also requires much fewer reconstructions in order to succeed, as compared to the L-curve method.
@@
Figure 10: Reconstructions of the Shepp-Logan brain phantom under radial sampling, with $NL_{\%}=15/100$ and $UR_{\%}=15\%$. The reconstruction obtained with the ALM is displayed in sub-figure A. The reconstruction obtained with the parameter that maximizes the mSSIM is displayed in sub-figure B. We observe that the two images are almost indistinguishable. Sub-figure C displays the ground-truth.
5 Conclusions
@@
ORCID
Gianluca Giacchi : $0$-0002-6809-1311
Isidoros Iakovidis : 0009-0005-2594-0028
Micah M. Murray : $0$-0002-7821-117X
Bastien Milani : $0$-0002-2410-2744
Benedetta Franceschiello : $0$-0003-0754-5081
Funding Information
This work was partially supported by Swiss National Science Foundation; grant 220433 to BF and the PON scholarship PONDOT1303154-3
Author contributions
G.G. conducted the analysis, developed the mathematical framework, implemented ALMA and provided a first draft of the article. I.I. developed the comparison with the L-curve. B.M. provided the code implementation and contributed to the algorithm design. B.F. conceived the idea of the work, and together with B.M. supervised its development. All authors contributed to the final version of the paper.
Financial disclosure
None reported.
Conflict of interest
The authors declare no potential conflict of interests.